Find the​ half-life of a radioactive​ element, which decays according to the function ​A(t)=A0e−0.0372t​, where t is the time in years.
Answer #1

SOLUTION;

Half life decay function is

A(t)=A_o e^{-0.0372t}

Where

A​​​​​​O = initial amount

A​​​​​(t)= remaining amount after t time

For , half life

So, A​​​​​​o = P ( supposed)

Then. A​​​​​(t)= P/2

Now,

A(t)=A_o e^{-0.0372t}

\frac{P}{2} =Pe^{-0.0372t}

\frac{1}{2} =e^{-0.0372t}

ln\frac{1}{2} =ln(e)^{-0.0372t}

ln\frac{1}{2} =-0.0372t

t= 18.6 years

The half life is = 18.6 years

Answer

18.6 years